# When LLMs Imagine People: A Human-Centered Persona Brainstorm Audit for Bias and Fairness in Creative Applications

### v1

@@ Abstract.
Biased outputs from Large Language Models (LLMs) can reinforce stereotypes and perpetuate inequities in real-world applications, making fairness auditing essential. We introduce the Persona Brainstorm Audit (PBA), a scalable and transparent auditing method for detecting bias through open-ended persona generation. Unlike existing methods that rely on fixed identity categories and static benchmarks, PBA uncovers biases across multiple social dimensions while supporting longitudinal tracking and mitigating data leakage risks. Applying PBA to 12 state-of-the-art LLMs, we compare bias severity across models, dimensions, and versions, uncover distinct patterns and lineage-specific variability, and trace how biases attenuate, persist, or resurface across successive generations. Robustness analyses show PBA remains stable under varying sample sizes, role-playing prompts, and debiasing prompts, establishing its reliability for fairness auditing in LLMs.
@@
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYccs: Human-centered computing Human computer interaction (HCI)$c$s: Human-centered computing Empirical studies in HCIccs: Computing methodologies Natural language processingccs: Social and professional topics Computing / technology policyccs: Social and professional topics User characteristics
@@ 1. Introduction
Large Language Models (LLMs) have advanced at an unprecedented pace, demonstrating remarkable capabilities in conversational-fluency, instruction-following, in-context learning, and reasoning (Achiam et al., 2023; Cao, 2024; Wang et al., 2025). These breakthroughs have driven widespread adoption across academia and industry, powering applications in information retrieval, customer service, education, healthcare, and creative domains (Arora et al., 2025; Pang et al., 2025). Their integration signals not only technical progress but also a profound shift in how knowledge work and creative practices are conducted (Chu et al., 2024; Amin et al., 2025). Recently, LLMs are increasingly applied in creative workflows, supporting tasks such as brainstorming (Fukumura and Ito, 2025), design (Lin and Martelaro, 2024), creative writing (Gero et al., 2023), persona generation (Amin et al., 2025; Sethi et al., 2025), and synthetic data generation (Cao, 2024). These applications promise efficiency and diversity of ideas, enabling designers and researchers to accelerate innovation.
@@
However, these opportunities come with critical challenges. LLMs often exhibit incomplete or outdated knowledge, hallucinations, and reliability issues (Cao, 2025). The rapid integration of LLMs into real-world applications can amplify existing bias and fairness challenges (Gallegos et al., 2024), undermining the reliability and robustness of Artificial Intelligence (AI) systems and exacerbating societal inequities. In creative contexts, these risks extend beyond factual accuracy: representational harms, stereotype reinforcement, and cultural homogenization can shape design outcomes and user experiences in subtle but consequential ways (Amin et al., 2025; Haque et al., 2025).
There is a rich body of research on bias detection in the area of Natural Language Processing (NLP) and language models. Early work examined bias in static word embeddings, using association tests such as the Word Embedding Association Test (WEAT) and its variants to quantify stereotypical associations (Bolukbasi et al., 2016; Caliskan et al., 2017). To extend these measures to contextualized embeddings, the Sentence Encoder Association Test (SEAT) was introduced to generate embeddings from semantically bleached and template-based sentences (May et al., 2019). While these approaches have been influential for probing internal representations, prior work shows that they have weak and inconsistent correlations with biases in downstream tasks and offer limited applicability to LLMs (Gallegos et al., 2024; Cabello et al., 2023; Delobelle et al., 2022). Subsequent work introduced masked token methods, which estimate token probabilities by masking words in a sentence and prompting a language model to predict the missing token. Although effective for examining contextual associations, these methods depend on rigid templates with limited semantic diversity and a narrow set of target words, which constrains their generalizability and robustness (Gallegos et al., 2024).
@@
Recent work has shifted toward LLM-specific bias detection, using metrics on generated text for tasks such as sentence completion (Dhamala et al., 2021) and question answering (Parrish et al., 2022). These approaches leverage prompt-based probing and generative benchmarks to assess stereotypes and fairness in generated outputs. However, they rely on user-defined datasets that suffer from ambiguity, limited coverage, and scalability issues (Gallegos et al., 2024; Davani et al., 2025). Moreover, these evaluations rarely reflect real-world usage, leaving creative downstream tasks such as persona generation and brainstorming largely unexamined. Unlike factual Question-Answering (QA), creative outputs influence downstream design decisions, making unexamined biases particularly consequential for inclusivity and fairness in real-world products. Prior work has examined bias and fairness in persona generation, including narrative style descriptive personas conditioned on specific domains or meta personas, through manual expert assessments and quantitative comparisons against ground truth profiles (Salminen et al., 2024; Li et al., 2025). These studies reveal persistent risks of stereotype propagation in creative applications. Despite these efforts, existing approaches remain fragmented and narrow in scope. Systematic, holistic, and scalable methods for examining bias and fairness in creative LLM outputs are still scarce, leaving practitioners without robust metrics to evaluate or mitigate these risks.
To address this gap, we introduce the Persona Brainstorm Audit (PBA), a scalable and interpretable auditing method for creative LLM use in persona generation. Our work is driven by a simple but pressing question: when LLMs imagine people, what patterns of bias emerge and how do they change over time? To unpack this, we explore three dimensions of inquiry. First, we ask whether leading LLMs differ in the severity and extent of bias when examined through PBA. Next, we investigate the shape of these biases by answering how do different LLMs exhibit distinct patterns of bias across identity dimensions. Finally, we examine how these patterns evolve across successive model releases, probing whether fairness improves, stagnates, or even regresses. The main contributions of this work are:
  • A novel auditing method: PBA is able to detect biases in LLM-generated personas without relying on fixed identity categories or static benchmarks.
  • A robust bias metric: We adapt and normalize Cramér’s V for multi-cardinality, multi-dimensional identity attributes, enabling fair and interpretable comparisons of bias strength across models and dimensions.
@@
  • A longitudinal, multi-model Analysis: We apply PBA to 12 state-of-the-art LLMs, uncovering nonlinear bias trajectories and domain-specific allocation patterns that challenge assumptions of monotonic fairness improvement.
@@
  • Actionable Implications: Design and governance recommendations for bias-aware creative AI, including continuous auditing, transparency practices, and sustainability considerations.
@@ 5.3. Impact of debiasing prompt
Name x Social Class 0.452 0.474 0.435 0.276 0.333 0.734
Name x Education 0.112 0.112 0.486 0.333 0.167 0.876
Name x Occupation 0.240 0.270 0.486 0.333 0.167 0.615
Name x Interest 0.540 0.561 0.543 0.467 0.333 0.965
Gender x Social Class 0.433 0.442 0.371 0.333 0.167 0.843
Gender x Education 0.450 0.352 0.543 0.467 0.167 0.869
Gender x Occupation 0.278 0.263 -0.029 -0.067 0.333 0.987
Gender x Interest 0.233 0.254 0.371 0.333 0.500 0.684
Ethnicity x Social Class 0.471 0.314 0.371 0.333 0.500 0.984
Ethnicity x Education 0.105 0.073 0.029 0.067 0.500 0.961
Ethnicity x Occupation 0.427 0.431 0.600 0.333 0.500 0.879
Ethnicity x Interest 0.828 0.844 0.714 0.467 0.333 0.997
Sexual Orientation x Social Class 0.103 0.087 0.314 0.200 0.000 0.837
Sexual Orientation x Education 0.174 0.076 0.174 0.138 0.167 0.880
Sexual Orientation x Occupation 0.662 0.606 0.771 0.600 0.333 0.815
Sexual Orientation x Interest 0.593 0.611 0.580 0.276 0.500 0.597
@@ 7. Conclusions
This work introduced the Persona Brainstorm Audit (PBA), a scalable and interpretable metric for auditing bias in LLM-generated personas. By moving beyond fixed identity categories, PBA surfaces multidimensional biases and enables longitudinal tracking across model generations. Applying PBA to 12 state-of-the-art LLMs, we show that bias trajectories are nonlinear and domain-specific, challenging assumptions that fairness improves with scale or alignment. Our findings highlight persistent structural patterns such as gendered occupational stereotypes and non-binary funneling, alongside unpredictable shifts across families and versions. These dynamics underscore the need for continuous auditing, model-specific mitigation strategies, and governance mechanisms that translate measurement into accountability. PBA’s robustness and extensibility make it a practical tool for fairness evaluation in creative AI applications. With generative AI being increasingly adopted in real-world applications, bias and fairness must be treated as an ongoing responsibility embedded in design, deployment, and policy.
Appendix A Glossary
This glossary provides definitions of acronyms and abbreviations used throughout the paper for quick reference:
  • AI – Artificial Intelligence
  • API – Application Programming Interface
  • ACL – Association for Computational Linguistics
  • CHI – Conference on Human Factors in Computing Systems
  • HCI – Human-Computer Interaction
  • ICC – Intraclass Correlation Coefficient
  • JSON – JavaScript Object Notation
  • LLM – Large Language Model
  • MTEB – Massive Text Embedding Benchmark
  • NLP – Natural Language Processing
  • PBA – Persona Brainstorm Audit
  • QA – Question Answering
  • SD – Severity Difference
  • SEAT – Sentence Encoder Association Test
  • UX – User Experience
  • WEAT – Word Embedding Association Test
Appendix B Supporting data
B.1. Cardinality Comparison Across Models
@@
B.2. Duplicate Profile Analysis
@@
B.3. Bias Visualization Across Dimensions
Analysis by model family shows distinct patterns of bias across generations of LLMs. For each subplot in Figure 6, the average bias score for each model series is displayed in the subplot title. OpenAI early frontier models (including GPT-3.5, GPT-4, and GPT-4o) have the lowest average bias scores. The Mistral series shows slightly higher averages. In contrast, the most recent OpenAI series, GPT-4.1 series and GPT-5 series, have the highest average bias scores. Within the GPT-4.1 series, the largest model GPT-4.1 has the lowest bias score. However, within the GPT-5 series, the smallest model GPT-5 nano has the lowest bias score. The Mistral family shows a similar pattern as GPT-5 series: the smallest model Ministral-3B has the lowest bias score. These findings reveal that bias trajectories in LLMs are neither linear nor uniform: lineage and scale interact in complex ways, producing structurally heterogeneous profiles where the magnitude and concentration of biases vary sharply across dimensions. This underscores that mitigating bias requires model-specific strategies rather than assumptions of steady progress across generations or model sizes.
B.4. Gender X Occupation analysis
@@ B.5. Zoom in analysis
B.6. Human Validation Study

### v2

@@ Abstract.
Large Language Models (LLMs) used in creative workflows can reinforce stereotypes and perpetuate inequities, making fairness auditing essential. Existing methods rely on constrained tasks and fixed benchmarks, leaving open-ended creative outputs unexamined. We introduce the Persona Brainstorm Audit (PBA), a scalable and easy to extend auditing method for bias detection across multiple intersecting identity and social roles in open-ended persona generation. PBA quantifies bias using degree-of-freedom-aware normalized Cramér’s V, producing interpretable severity labels that enable fair comparison across models and dimensions. Applying PBA to 12 LLMs (120,000 personas, 16 bias dimensions), we find that bias evolves nonlinearly across model generations: larger and newer models are not consistently fairer, and biases that initially decrease can resurface in later releases. Intersectional analysis reveals disparities hidden by single-axis metrics, where dimensions appearing fair individually can exhibit high bias in combination. Robustness analyses show PBA remains stable under varying sample sizes, role-playing prompts, and debiasing prompts, establishing its reliability for fairness auditing in LLMs.
@@
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NY
@@ 1. Introduction
Large Language Models (LLMs) have advanced rapidly, driving widespread adoption across academia and industry (Cao, 2024; Wang et al., 2025; Pang et al., 2025). In particular, LLMs are increasingly integrated into creative tasks such as brainstorming (Fukumura and Ito, 2025), design (Lin and Martelaro, 2024), persona generation (Amin et al., 2025), and synthetic data generation (Cao, 2024), promising efficiency and diversity while reshaping creative practices (Chu et al., 2024; Amin et al., 2025). However, LLMs can exhibit reliability issues and hallucinations (Cao, 2025), and their rapid integration into real-world applications can amplify bias and fairness challenges, undermining AI robustness and exacerbating societal inequities (Gallegos et al., 2024). In creative contexts, these risks manifest as representational harms, stereotype reinforcement, and cultural homogenization that shape downstream outcomes in consequential ways (Amin et al., 2025; Haque et al., 2025).
@@
Bias detection has progressed from embedding-based methods (Bolukbasi et al., 2016; Caliskan et al., 2017) to masked token approaches, but these show weak correlations with downstream tasks, limited applicability to LLMs (Gallegos et al., 2024; Cabello et al., 2023), and rely on rigid templates with narrow semantic diversity (Gallegos et al., 2024). Recent work has shifted toward generation-based bias detection using sentence completion (Dhamala et al., 2021) and question answering (Parrish et al., 2022), but these approaches rely on constrained datasets with limited coverage and scalability (Gallegos et al., 2024; Davani et al., 2025) and rarely reflect real-world creative usage. Unlike factual QA, creative outputs such as persona generation directly influence downstream decisions, making unexamined biases particularly consequential; prior work on persona generation has examined bias through manual expert assessments and ground truth comparisons (Li et al., 2025), revealing persistent stereotype risks. Yet existing approaches remain fragmented: systematic, scalable, and interpretable methods for bias detection in creative LLM outputs are still scarce.
To address this gap, we introduce the Persona Brainstorm Audit (PBA), a scalable and interpretable method for bias detection in creative LLM tasks. Specifically, we investigate what patterns of bias emerge when LLMs perform creative tasks, how these patterns vary across identity dimensions, and how they evolve across model generations. The main contributions of this work are:
  • A novel auditing method: PBA detects systematic bias in LLM-generated personas across multiple intersecting identity dimensions without relying on fixed categories, constrained tasks, or external classifiers.
  • An interpretable, cross-comparable bias metric: We normalize Cramér’s V with degree-of-freedom-aware thresholds, producing human-readable severity labels that enable fair comparison across models and dimensions.
@@
  • A large-scale longitudinal analysis: We apply PBA to 12 LLMs (120,000 personas), uncovering nonlinear bias trajectories that challenge assumptions of monotonic fairness improvement.
@@ 5.3. Impact of debiasing prompt
Name x Social Class 0.452 0.474 0.435 0.276 0.333 0.741
Name x Education 0.112 0.112 0.486 0.333 0.167 0.823
Name x Occupation 0.240 0.270 0.486 0.333 0.167 0.633
Name x Interest 0.540 0.561 0.543 0.467 0.333 0.255
Gender x Social Class 0.433 0.442 0.371 0.333 0.167 0.791
Gender x Education 0.450 0.352 0.543 0.467 0.167 0.949
Gender x Occupation 0.278 0.263 -0.029 -0.067 0.333 0.861
Gender x Interest 0.233 0.254 0.371 0.333 0.500 0.288
Ethnicity x Social Class 0.471 0.314 0.371 0.333 0.500 0.975
Ethnicity x Education 0.105 0.073 0.029 0.067 0.500 0.949
Ethnicity x Occupation 0.427 0.431 0.600 0.333 0.500 0.808
Ethnicity x Interest 0.828 0.844 0.714 0.467 0.333 0.709
Sexual Orientation x Social Class 0.103 0.087 0.314 0.200 0.000 0.902
Sexual Orientation x Education 0.174 0.076 0.174 0.138 0.167 0.988
Sexual Orientation x Occupation 0.662 0.606 0.771 0.600 0.333 0.918
Sexual Orientation x Interest 0.593 0.611 0.580 0.276 0.500 0.759
Mean 0.320 0.319 -0.200 -0.200 0.167 0.822
@@ 7. Conclusions
This work introduced the Persona Brainstorm Audit (PBA), a scalable method for auditing bias in open-ended LLM-generated personas using degree-of-freedom-aware normalized Cramér’s V, producing bounded severity labels that enable fair comparison across models and dimensions without classifiers or fixed-answer templates. Applying PBA to 12 LLMs across 120,000 personas and 16 bias dimensions, we find that bias trajectories are nonlinear and architecture-dependent: early frontier models (GPT-4o, GPT-4) are among the least biased, while more recent and larger models (GPT-4.1 mini, GPT-5) rank among the most biased. Dimension-level and intersectional analyses reveal that identities appearing fair in isolation can produce compounding disparities in combination, and that stereotype patterns such as occupational clustering by sexual orientation are persistent and model-family-specific. Robustness analyses confirm PBA’s stability under varying sample sizes and prompt conditions, though prompt-based debiasing alone proves insufficient for meaningful bias reduction. A current limitation is PBA’s focus on personas in English; future work will extend to multilingual settings and investigate causal mechanisms behind nonlinear bias patterns.
Appendix A Supporting data
A.1. Cardinality Comparison Across Models
@@
A.2. Duplicate Profile Analysis
@@
A.3. Gender X Occupation analysis
@@ A.4. Zoom in analysis
A.5. Human Validation Study